Glariel:

From the top of the tower 63.2 feet high, the angle of depression of the two objects situated in the same horizontal line with the base of the building, and on the same side of the tower are 31° 16’ and 48° 28’ respectively. Find the distance between the two objects

jhonyy9:

 find distance AB

jhonyy9:

@AZ any idea ? opinion pls - ty

AZ:

Once again, a very helpful drawing! Our angles are 31 degrees and 16 minutes and 48 degrees and 28 minutes To convert it to only degrees, we have to divide the minutes number by 60 (since there are 60 minutes in 1 degree. So for example, if we had 25 degrees and 15 minutes \(\dfrac{\text{15 minutes}}{1} \times \dfrac{\text{1 degree}}{\text{60 minutes}}\) and we get that 15' or 15 minutes is 0.25 degrees so 25 degrees and 15 minutes is 15.25 degrees or 15.25°

AZ:

So for the first step, can you convert 31 degrees and 16 minutes and 48 degrees and 28 minutes into just degrees? Just divide the minutes part by 60 and that will be the decimal that comes after the degrees :)

AZ:

The second step that you have to do is you a trigonometric function to calculate the length of BC and then of AC We can then do AC - BC to calculate the length from A to B which is the distance between the two objects and it's what the question wants us to find

AZ:

Responding to this post to post a correct drawing since it's angle of DEPRESSION

1 attachment
AZ:

We have an angle. We have the side adjacent to it. And we're looking to find the distance of the opposite side. So which trigonometric function can we use?

1 attachment